# Estimating Correctness Without Oracles in LLM-Based Code Generation

### v1

@@ 5.1 Fixed Sampling Budget for $\mathsf{Coder}(d)$
In theory, pointwise incoherence can be estimated efficiently using simple Monte Carlo procedures. The estimator defined by Equation (8) is easy to implement, parallelizable, and statistically robust. As shown in Theorem 4.2 and Theorem 4.3, both incoherence estimation and detection admit PAC guarantees with low sample complexity.
@@
While this approximation introduces some additional variance, it is highly effective in practice. It amortizes LLM sampling costs across many evaluations, enabling scalable incoherence and error estimation over large benchmark suites and ablations. In our experiments, we find that a larger choice of $m$ consistently yields more reliable incoherence and error estimates across diverse tasks.
@@ 5.2 Test Input Generation to Implement $\mathsf{Gen}$
Automatic software test input generation is a well-studied problem in the software engineering community. Cast as a constraint satisfaction problem, we can use symbolic execution to generate inputs that exercise the different paths of a program (King 1976) or that reveal a difference between two program versions (Böhme, Oliveira, and Roychoudhury 2013). Cast as an optimization problem, we can use heuristic search to generate inputs that maximize code coverage (Ferguson and Korel 1996).
For our purposes, we propose to use fuzzing, an approach that mutates a set of user-provided or auto-generated seed inputs to generate new inputs. Today, fuzzing is the most successful and most widely-deployed automatic testing technique in practice (Böhme, Cadar, and Roychoudhury 2021). Like random test input generation, fuzzing is amenable to statistical guarantees, e.g., to quantify the probability of finding a bug with the next generated input in an ongoing testing campaign that has found no bugs (Böhme 2019, 2018; Böhme, Liyanage, and Wüstholz 2021).
In fact, fuzzing has recently been proposed specifically to improve the soundness of the evaluation of LLM-based code generators on the HumanEval and MBPP benchmarks (Liu et al. 2023), where pass@1 (i.e., mean error across all benchmark tasks) was traditionally computed using five test inputs per task (Chen et al. 2021). The technique EvalPlus constructs the input distributions $\mathsf{Gen}$ in two stages:
@@
We observe that Stage 1 might introduce a bias where an LLM’s generated code might appear to perform better on inputs generated by the same LLM, compared to inputs generated by another LLM. Hence, in our experiments, to provide a fair evaluation of all considered LLMs, we mitigate that potential bias by using the benchmark-provided test inputs as seed inputs. In practice, in the absence of existing test inputs, we suggest using the original method or discovering the seed corpus using greybox fuzzing (at the cost of testing efficiency) (Zalewski 2014).
@@ 6.2 Models and Datasets
Table 1: Large Language Models used in our experiments.
@@
Models. Table 1 shows the large language models (LLMs) used in our experiments. At the time of writing, these 16 LLMs represent the most successful LLMs for code generation according to several popular leaderboards (Leaderboards 2025). They also represent the current portfolio of the most popular LLM vendors: Anthropic, DeepSeek, Google, Meta, Mistral, and OpenAI. By default, we chose a temperature of 0.6, a value commonly used in prior work on code generation (Li et al. 2024; DeepSeek-AI et al. 2025). We vary the temperature parameter in the ablation study (RQ3).
Datasets. We evaluate our measures of incoherence and error using the 16 LLMs on two (2) popular code generation benchmarks: HumanEval (Ji et al. 2025) and MBPP (Mostly Basic Python Problems) (Hu et al. 2025). HumanEval is a human-written benchmark published by OpenAI in 2021, consisting of 164 programming tasks. MBPP is a crowd-sourced benchmark published by Google in 2022. We used the author-sanitized version of MBPP containing 426 hand-verified programming tasks. For every task, they offer
@@ 6.3 Variables and Measures
Given a code generator $\mathsf{Coder}$ and input generator $\mathsf{Gen}$, programming task $d$, a query budget $m$ and a testing budget $n$, the empirical error ^$\hat{\operatorname{\mathcal{E}}}(d,m,n)$ on $d$ as estimator of the pointwise error is computed as
^$\hat{\operatorname{\mathcal{E}}}(d,m,n)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(\llbracket\pi^{d}_{y_{i}}\rrbracket(x^{d}_{i})\neq f^{\ast}_{d}(x^{d}_{i}))$ (9)
and the empirical incoherence ^$\hat{\operatorname{\mathcal{I}}}(d,m,n)$ on $d$ as estimator of the pointwise incoherence is computed as
^$\hat{\operatorname{\mathcal{I}}}(d,m,n)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(\llbracket\pi^{d}_{y_{i}}\rrbracket(x_{i}^{d})\neq\llbracket\pi^{d}_{y^{\prime}_{i}}\rrbracket(x_{i}^{d}))$ (10)
@@
Given programming tasks $S$, we compute the mean empirical error and mean empirical incoherence as
¯$\bar{\operatorname{\mathcal{E}}}(S,m,n)$ $=\frac{1}{|S|}\sum_{d\in S}\hat{\operatorname{\mathcal{E}}}(d,m,n)$ (11)
¯$\bar{\operatorname{\mathcal{I}}}(S,m,n)$ $=\frac{1}{|S|}\sum_{d\in S}\hat{\operatorname{\mathcal{I}}}(d,m,n)$ (12)
We can now write the empirical pass@1 score in terms of the mean empirical error (cf. Eq. (4)):
$1-\frac{1}{|S|}\sum_{d\in S}\mathbb{I}\left(0\neq\sum_{j=1}^{n}\mathbb{I}\left(\llbracket\pi^{d}_{1}\rrbracket(x^{d}_{j})\neq f^{\ast}_{d}(x^{d}_{j})\right)\right)$ (13)
$=1-\bar{\operatorname{\mathcal{E}}}(S,1,n)$ (14)
The detection rate is the proportion of tasks with non- zero emp. error that have a non-zero emp. incoherence.
$\frac{1}{|S_{x}|}\sum_{d\in S_{x}}\mathbb{I}(0\neq\hat{\operatorname{\mathcal{I}}}(d,m,n))$ (15)
where $S_{x}=\{d\ |\ d\in S\wedge\hat{\operatorname{\mathcal{E}}}(d,m,n)\neq 0\}$.
The undetected mean empirical error is the mean emp. error of tasks with zero emp. incoherence.
$\frac{1}{|S_{u}|}\sum_{d\in S_{u}}\hat{\operatorname{\mathcal{E}}}(d,m,n)$ (16)
where $S_{u}=\{d\ |\ d\in S\wedge\hat{\operatorname{\mathcal{I}}}(d,m,n)=0\}$.
We measure the strength of the relationship between two random variables, i.e., empirical incoherence and error, using Spearman’s rank correlation coefficient $\rho$.
We measure the agreement on ranking when sorting the performance of LLMs measured by the proportion of programming tasks (a) with zero mean empirical error versus (b) with zero mean empirical incoherence, also using Spearman’s rank correlation coefficient.
Figure 2: Workflow of our implementation DiffTrust.
@@ 6.4 Implementation
For every programming task, once for the empirical error and once for the empirical incoherence, DiffTrust uses the input generator to generate $n$ inputs for the generated candidate functions using the task-provided seed inputs. The test generator implements the budget-constrained $\mathsf{Coder}_{m}$ presented in Section 5.1 and the type-aware mutation-based fuzzing method $\mathsf{Gen}$ introduced in EvalPlus (Liu et al. 2023) and discussed in Section 5.2. We provide a list of supported mutations in the appendix (Table 6). To ensure robustness, all executions, whether for compilation, incoherence estimation, or error estimation, are subject to a 60-second timeout. The empirical error is computed using the task-provided ground-truth (GT) function $f^{*}_{d}$. By default, we generate $n=1000$ test inputs. We vary $n\in\{100,1000,2000,5000,10000\}$ in RQ3.
https://github.com/mpi-softsec/difftrust
6.5 Infrastructure
@@ 7 Empirical Results
Human Eval Code 0.0763 0.0295 0.7171 0.7188 0.0417
@@
Results. A non-zero incoherence effectively detects a non-zero error without access to a ground truth implementation. The mean detection rate across all 16 LLMs for MBPP and HumanEval are 69% and 66%, respectively. There does not seem to be a substantial difference in detection rate between the code-generation specific LLM (Gemini 2.5 Pro) and the general-purpose or the small LLM (GPT-4o, Ministral 8b). In cases where the incoherence is zero, the mean error is also substantially lower. Concretely, the mean error reduces from 30% to 19% for MBPP and from 11% to 5% for HumanEval (“Undetected Mean Error”). For the small LLM, the mean error is generally higher than for the other LLMs, but the percentage decrease when incoherence is zero is similar.
Figure 3 best illustrates the relationship between error and incoherence for GPT-4o (called general in the table). We can clearly see the consequence of the inequality as shown in Theorem 4.1. The error is usually greater than incoherence for a programming task. There are some tasks (on the left of each plot) where incoherence is zero but the error is non-zero. For the average LLM, we find a moderate correlation between error and incoherence (0.56 for MBPP; 0.69 for HumanEval). In Table 2, for the three LLMs evaluated on HumanEval, we even find a strong correlation (gt. 0.7).
@@
If the rankings agree, we can reliably substitute one measure for the other. We could remove the requirement to provide painstakingly manually-written ground-truth implementations for every programming task when constructing new code generation benchmarks. We could mitigate critical threats to validity in benchmarking of new LLM-based code generation systems, such as overfitting or data leakage.
Results. We observe a very strong agreement on the rankings, which is quantified by a Spearman correlation of 0.92 and 0.94 for MBPP and HumanEval, respectively, at a significance level $p<0.0001$. The rankings are close to the diagonal. We can reliable substitute one measure for the other.
@@
Query budget $m$. The detection rate increases as the query budget increases for both benchmarks. For instance, when changing $m$ from 10 to 50, we see the detection rate increase by 14–19% from 0.7 to about 0.83 for HumanEval and from 0.72 to 0.82 for MBPP. The increase in detection rate comes at a substantial monetary cost. When changing $m$ from 10 to 50, our expenses increased by more than fivefold, e.g., from \$9 to \$44 for MBPP. For HumanEval, we actually observe a slightly higher detection rate (0.83) at $m=25$, which we explain by the randomness of the sampling and test generation process. It might also indicate that the detection rate starts to saturate for larger values of $m$ (which we determined as uneconomical for us to test). Another interesting observation is that just sampling a second program ($m=2$) already gives us a 0.33 to 0.4 detection rate.
Temperature $t$. Detection rate increases as the model’s temperature increases for both benchmarks. For instance, when changing $t$ from 0.2 to 1.0, we see the detection rate increase by 36–50% from 0.54 to 0.81 for MBPP and from 0.56 to 0.76 for HumanEval. A high temperature induces a high output diversity, which seems to increase the LLM’s incoherence on that programming task, which serves us well in the detection of errors.
Test inputs $n$. Detection rate increases as the number of generated test inputs increases. However, compared to the other hyperparameters, a substantial increase in the number of generated test inputs induces only a relatively small increase in detection rate.
@@ 8 Threats to Validity
The second threat is to the internal validity, i.e., the extent to which the presented evidence supports our claims about cause and effect within the context of our study. In the benchmarks, the task description may be ambiguous or the ground-truth implementation incorrect (Siddiq et al. 2024). We use popular well-scrutinized benchmarks. From MBPP, we chose the best-quality, hand-curated set of tasks. DiffTrust may contain bugs itself, but we release all our scripts and data for the community to scrutinize.
@@ 9 Perspective
In this paper, we discuss how to formally estimate the correctness of a program generated to solve a programming task when there is no automated mechanism to decide whether a program is correct or not (e.g., a formal specification or a ground-truth implementation). We model the generated program as a random variable drawn from an unknown distribution induced by the coder (e.g., the LLM). This opens the door for a probabilistic notion of correctness. Our measure, incoherence, formalizes our observation that, if two random programs for the same programming task disagree on the output for an input, at least one of them must be incorrect. We formally demonstrate how the coder’s incoherence on a task provides a provable lower bound on the coder’s error on that task and empirically observe that a non-zero incoherence detects more than two-thirds of the incorrect programs (where detection rate increases with the number of generated candidates and test inputs). Since incoherence depends solely on observable outputs and not on model internals, it can be applied broadly across LLMs of any kind and even to probabilistic systems like LLM-agent architectures.
@@
We believe this opens up many possibilities, both when using LLMs to write programs, as well as for evaluating the code generation capabilities of multiple LLMs (or agents) simultaneously. In particular, the very strong agreement between the ranking of 16 LLMs in terms of our ground-truth-less incoherence-based measure versus the ranking in terms of the existing ground-truth-based pass@1 opens up exciting possibilities. The existing process of curating large coding benchmarks with correct human-generated ground-truth implementations is labour-intensive, error-prone, and subject to future data leakage issues (Ramos et al. 2025). Incoherence paves the way for evaluations on a substantially larger scale, basically on a stream of programming tasks.
Acknowledgements
@@
This research is partially funded by a research scholarship from ENS Paris-Saclay for an internship at the MPI-SP Software Security group. This research is also partially funded by the European Union. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work is supported by an ERC grant (Project AT_SCALE, 101179366).
Appendix A Proof of Pointwise Incoherence Inequality

### v2

@@ 5.1 Fixed Sampling Budget for $\mathsf{Coder}(d)$
In theory, pointwise incoherence can be estimated efficiently. The estimator defined by Equation (8) is easy to implement, parallelizable, and statistically robust. Voth incoherence estimation and detection admit PAC guarantees with low sample complexity (cf. Thm. 4.2, Thm. 4.3).
@@
While this approximation introduces some additional variance, it is highly effective in practice. It amortizes LLM sampling costs across many evaluations, enabling scalable incoherence and error estimation. Experimentally, we find that a larger $m$ consistently yields more reliable estimates.
@@ 5.2 Test Input Generation to Implement $\mathsf{Gen}$
Automatic software test input generation is a well-studied problem in the software engineering community. Cast as a constraint satisfaction problem, we can use symbolic execution to generate inputs that exercise the different paths of a program (king) or that reveal a difference between two program versions (prv). Cast as an optimization problem, we can use heuristic search to generate inputs that maximize code coverage (korel).
For our purposes, we propose to use fuzzing, an approach that mutates a set of user-provided or auto-generated seed inputs to generate new inputs. Today, fuzzing is the most successful and most widely-deployed automatic testing technique in practice (fuzzing). Like random test input generation, fuzzing is amenable to statistical guarantees, e.g., to quantify the probability of finding a bug with the next generated input in an ongoing testing campaign that has found no bugs (assurance; stads; residualrisk; ICSE26-dependency; ICLR25-unseen; reachability).
In fact, fuzzing has recently been proposed specifically to improve the soundness of the evaluation of LLM-based code generators on the HumanEval and MBPP benchmarks (EvalPlus), where pass@1 (i.e., mean error across all benchmark tasks) was traditionally computed using five test inputs per task (chen2021evaluating). The technique EvalPlus constructs the input distributions $\mathsf{Gen}$ in two stages:
@@
We observe that Stage 1 might introduce a bias where an LLM’s generated code might appear to perform better on inputs generated by the same LLM, compared to inputs generated by another LLM. Hence, in our experiments, to provide a fair evaluation of all considered LLMs, we mitigate that potential bias by using the benchmark-provided test inputs as seed inputs. In practice, in the absence of existing test inputs, we suggest using the original method or discovering the seed corpus using greybox fuzzings (thinair).
@@ 6.2 Models and Datasets
Table 1: Large Language Models used in our experiments.
Models. Table 1 shows the large language models (LLMs) used in our experiments. At the time of writing, these 16 LLMs represent the most successful LLMs for code generation according to several popular leaderboards (leaderboards). They also represent the current portfolio of the most popular LLM vendors: Anthropic, DeepSeek, Google, Meta, Mistral, and OpenAI. By default, we chose a temperature of 0.6, a value commonly used in prior work on code generation (li2024honestcoder; deepseek). We vary the temperature parameter in the ablation study (RQ3).
Datasets. We evaluate our measures of incoherence and error using the 16 LLMs on two (2) popular code generation benchmarks: HumanEval (humaneval) and MBPP (Mostly Basic Python Problems) (MBPP). HumanEval is a human-written benchmark published by OpenAI in 2021, consisting of 164 programming tasks. MBPP is a crowd-sourced benchmark published by Google in 2022. We used the author-sanitized version of MBPP containing 426 hand-verified programming tasks. For every task, they offer
@@ 6.3 Variables and Measures
Given a code generator $\mathsf{Coder}$ and input generator $\mathsf{Gen}$, programming task $d$, a query budget $m$ and a testing budget $n$, the empirical error $\hat{\operatorname{\mathcal{E}}}(d,m,n)$ on $d$ as estimator of the pointwise error is computed as
$\hat{\operatorname{\mathcal{E}}}(d,m,n)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\bigl(\llbracket\pi^{d}_{y_{i}}\rrbracket(x^{d}_{i})\neq f^{\ast}_{d}(x^{d}_{i})\bigr)$ (9)
and the empirical incoherence $\hat{\operatorname{\mathcal{I}}}(d,m,n)$ on $d$ as estimator of the pointwise incoherence is computed as
$\hat{\operatorname{\mathcal{I}}}(d,m,n)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\bigl(\llbracket\pi^{d}_{y_{i}}\rrbracket(x_{i}^{d})\neq\llbracket\pi^{d}_{y^{\prime}_{i}}\rrbracket(x_{i}^{d})\bigr)$ (10)
@@
Given the set of programming tasks $S$, we can now write the empirical pass@1 score in terms of the empirical error (cf. Eq. (4)):
$1-\frac{1}{|S|}\sum_{d\in S}\mathbb{I}\bigl(0\neq\hat{\operatorname{\mathcal{E}}}(d,1,n)\bigr)$ (11)
The mean empirical error $\bar{\operatorname{\mathcal{E}}}(S,m,n)$ is $\sum_{d\in S}\frac{\hat{\operatorname{\mathcal{E}}}(d,m,n)}{|S|}$ while the mean empirical incoherence $\bar{\operatorname{\mathcal{I}}}(S,m,n)$ is computed as $\sum_{d\in S}\frac{\hat{\operatorname{\mathcal{I}}}(d,m,n)}{|S|}$.
The detection rate is the proportion of tasks with non- zero emp. error that have a non-zero empirical incoherence, i.e., $\frac{1}{|S_{x}|}\sum_{d\in S_{x}}\mathbb{I}\bigl(0\neq\hat{\operatorname{\mathcal{I}}}(d,m,n)\bigr)$ where $S_{x}=\{d\ |\ d\in S\wedge\hat{\operatorname{\mathcal{E}}}(d,m,n)\neq 0\}$.
The undetected mean empirical error is the mean empirical error of tasks with zero empirical incoherence, i.e., $\frac{1}{|S_{u}|}\sum_{d\in S_{u}}\hat{\operatorname{\mathcal{E}}}(d,m,n)$ where $S_{u}=\{d\ |\ d\in S\wedge\hat{\operatorname{\mathcal{I}}}(d,m,n)=0\}$.
We measure the strength of the relationship between two random variables, i.e., empirical incoherence and error, using Spearman’s rank correlation coefficient $\rho$. We measure the agreement on ranking when sorting the performance of LLMs measured by the proportion of programming tasks (a) with zero mean empirical error versus (b) with zero mean empirical incoherence, also using Spearman’s rank correlation coefficient.
@@ 6.4 Implementation
For each task, we generate $m$ candidate functions per coder (default $m=10$), using vendor APIs. Inputs are generated via mutation-based fuzzing ($n=1000$ by default). We estimate both error (w.r.t. ground truth $f^{*}_{d}$) and incoherence (between candidates), with all executions sandboxed (60s timeout). Experiments ran on AMD EPYC 7713P CPU (128 threads), 251 GB RAM. More details in Appendix A.
@@ 7 Empirical Results
Human
Eval
Code 0.0763 0.0295 0.7171 0.7188 0.0417
@@
Results. A non-zero incoherence effectively detects a non-zero error without access to a ground truth implementation. The mean detection rate across all 16 LLMs for MBPP and HumanEval are 69% and 66%, respectively. There does not seem to be a substantial difference in detection rate between the code-generation specific LLM (Gemini 2.5 Pro) and the general-purpose or the small LLM (GPT-4o, Mistral 8b). In cases where the incoherence is zero, the mean error is also substantially lower. Concretely, the mean error reduces from 30% to 19% for MBPP and from 11% to 5% for HumanEval (“Undetected Mean Error”). For the small LLM, the mean error is generally higher than for the other LLMs, but the percentage decrease when incoherence is zero is similar.
Figure 2 best illustrates the relationship between error and incoherence. We can clearly see the consequence of the inequality in Theorem 4.1. Error is usually greater than incoherence for a programming task. There are some tasks (on the left of each plot) where incoherence is zero but the error is non-zero. For the average LLM, we find a moderate correlation between error and incoherence (0.56 for MBPP; 0.69 for HumanEval). In Table 2, for the three LLMs evaluated on HumanEval, we even find a strong correlation (gt. 0.7).
@@
If the incoherence- and error-based rankings of LLMs agree, we can reliably substitute one measure for the other. We could remove the requirement to provide painstakingly manually-written ground-truth implementations for every programming task when constructing new code generation benchmarks. We could mitigate critical threats to validity in benchmarking of new LLM-based code generation systems, such as overfitting or data leakage.
@@
Results. We observe a very strong agreement on the rankings. Concretely, $\rho=0.92$ and $\rho=0.94$ for MBPP and HumanEval, respectively, at a significance level $p<0.0001$. The rankings are close to the diagonal.
@@
Query budget $m$. The detection rate increases with $m$. For instance, when changing $m$ from 10 to 50, the detection rate increases by 14–19% from 0.7 to about 0.83 for HumanEval and from 0.72 to 0.82 for MBPP. The advantage comes at a substantial monetary cost. When changing $m$ from 10 to 50, our expenses increased by more than fivefold, e.g., from \$9 to \$44 for MBPP. For HumanEval, we actually observe a slightly higher detection rate (0.83) at $m=25$, which we explain by the randomness of the sampling and test generation process. It might also indicate that the detection rate starts to saturate for larger values of $m$ (which we determined as uneconomical for us to test). Another interesting observation is that just sampling a second program ($m=2$) already gives us a 0.33 to 0.4 detection rate.
Temperature $t$. Detection rate increases with $t$. For instance, when changing $t$ from 0.2 to 1.0, we see detection rate increase by 36–50% from 0.54 to 0.81 for MBPP and from 0.56 to 0.76 for HumanEval. A high temperature induces a high output diversity, which seems to increase the LLM’s incoherence, which serves us well in error detection.
Test inputs $n$. Detection rate increases with $n$. However, compared to the other hyperparameters, a substantial increase in the number of generated test inputs induces only a relatively small increase in detection rate.
@@ 8 Threats to Validity
The second threat is to the internal validity, i.e., the extent to which the presented evidence supports our claims about cause and effect within the context of our study. In the benchmarks, the task description may be ambiguous or the ground-truth implementation incorrect (qualityassessment). We use popular well-scrutinized benchmarks. From MBPP, we chose the best-quality, hand-curated set of tasks. DiffTrustmay contain bugs itself, but we release all our scripts and data for the community to scrutinize.
@@ 9 Perspective
In this paper, we discuss the formal estimation of the correctness of an LLM-generated program when there is no automated mechanism to decide whether a program is correct or not (e.g., a formal specification or a ground-truth implementation). We model the generated program as a random variable drawn from an unknown distribution induced by the coder (e.g., the LLM). This opens the door for a probabilistic notion of correctness. Our measure, incoherence, formalizes the observation that, if two random programs for the same task disagree on the output for an input, at least one must be incorrect. We formally demonstrate how the coder’s error on a task has a lower bound in terms of the coder’s incoherence on that task and empirically observe that a non-zero incoherence detects more than two-thirds of the incorrect programs. Since incoherence does not depend on model internals, it can be applied broadly across LLMs of any kind and even to probabilistic systems like LLM-agents.
@@
An incoherence-based evaluation of the code generation capabilities of multiple LLMs also addresses several open challenges of the traditional ground-truth-based pass@1 evaluation. The existing process of curating large coding benchmarks with correct human-generated ground-truth implementations is labour-intensive, error-prone, and subject to future data leakage issues (memorizing). Incoherence paves the way for evaluations on a substantially larger scale, basically on a stream of programming tasks.
Acknowledgments
@@
We thank the anonymous reviewers for their constructive feedback and for helping us improve this paper. This research is partially funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work is supported by ERC grant (Project AT_SCALE, 101179366). This research is also partially funded by the ENS Paris-Saclay who graciously supported the internship of the first author at the MPI for Security and Privacy.
Appendix A Implementation Details
For every programming task, once for the empirical error and once for the empirical incoherence, DiffTrustuses the input generator to generate $n$ inputs for the generated candidate functions using the task-provided seed inputs. The test generator implements the budget-constrained $\mathsf{Coder}_{m}$ presented in Section 5.1 and the type-aware mutation-based fuzzing method $\mathsf{Gen}$ introduced in EvalPlus (EvalPlus) and discussed in Section 5.2. We provide a list of supported mutations in the appendix (Table 6). We note that non-zero incoherence implies non-zero error for all $\mathsf{Gen}$ (Thm. 4.1 and Eqn. (7)). We also note that point-wise error [Eqn. (6)] and incoherence [Eqn. (7)] are defined w.r.t. the same distribution induced by $\mathsf{Gen}$. To ensure robustness, all executions, whether for compilation, incoherence estimation, or error estimation, are subject to a 60-second timeout. The empirical error is computed using the task-provided ground-truth (GT) function $f^{*}_{d}$. By default, we generate $n=1000$ test inputs. We vary $n\in\{100,1000,2000,5000,10000\}$ in RQ3.
Figure 3: Workflow of our implementation DiffTrust.
https://github.com/mpi-softsec/difftrust
A.1 Infrastructure
Appendix B Proof of Pointwise Incoherence Inequality
For readers who prefer a fully explicit probability-theoretic formalization, we describe here the underlying probability space used implicitly in the main text.
We consider a probability space $(\Omega,\mathcal{F},\mathbb{P})$, where $\Omega$ is the sample space, $\mathcal{F}$ the $\sigma$-algebra of events, and $\mathbb{P}$ the probability measure. Random variables are measurable functions from $\Omega$ to their codomain. For a mathematical expression $expr[X_{1},\dots,X_{n}]$ involving random variables, we write
$\{expr\}:=\{\omega\in\Omega\mid expr[X_{1}(\omega),\dots,X_{n}(\omega)]\text{ holds}\}.$